OpenStudy (anonymous):

f(x)=x-1/x squared+1 g(x)=x+10/x+13 what is (f o g)(x)?

OpenStudy (anonymous):

\[\frac{x-1}{x^2-1}\]?

OpenStudy (anonymous):

or maybe \[\frac{x-1}{x^2+1}\]?

OpenStudy (anonymous):

and \[g(x)=\frac{x+10}{x+13}\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

if so then \[f\circ g(x)= f(g(x)=f(\frac{x+10}{x+13})\] \[=\frac{\frac{x+10}{x+13}-1}{\frac{x+10}{x+13}+1}\]

OpenStudy (anonymous):

not that either